Crystal Glassware Company issues $1,043,000 of it's 13​%, 10-year bonds at 99 on February​ 28, 2019. The bonds pay interest on February 28 and August 31. Assume that Crystal uses the straight-line method for amortization. What net amount will be reported for the bonds on the August​ 31, 2019 balance​ sheet?
Answer #1

Answer

  • Net amount will be reported for the bonds on the August​ 31, 2019 balance​ sheet
    = $ 1,033,092 or $ 1,033,091.50
  • Working

A

Face Value

$1,043,000.00

B = A x 99/100

Issue Price

$1,032,570.00

C = A - B

Discount on Bonds Payable

$10,430.00

D

Life (years)

                                10

E = D x 2

No of semi annual interest payments

                                20

F = C/E

Straight Line amortisation

$521.50

G = C - F

Unamortised discount balance on Aug 31, 2019 (after #1 semi annual payment)

$9,908.50

H = A - G

Net Amount at which the Bonds will be reported at

$1,033,091.50

Answer